Calculate the molar solubility of AgCl in a 1.00−L solution containing 10.1 g of dissolved CaCl2. The Ksp for silver chloride is 1.6 × 10−10.
Answer #1

concentration of Cacl2 = (w/Mwt)*1/V

                       = (10.1/111)*(1/1)

                       = 0.091 M

[Cl-] = 2*0.091 = 0.182 M

Ksp of AgCl = [Ag+][Cl-]

1.6*10^-10 = s*(0.182)

s = solubility of AgCl = 8.8*10^-10 M